OpenStudy (safia21):

Two airplanes start at the same time from airports 500 km apart. Each one flies with an airspeed of 200km/h directly toward the other airport. But one reaches its airport half an hour before the other plane reaches the other airport. How fast is the wind blowing?

OpenStudy (amistre64):

i take it the planes are flying away from each other

OpenStudy (anonymous):

This feels like an implicit differentiation problem

OpenStudy (safia21):

its not its a rate problem

OpenStudy (amistre64):

s = d/t ; t = d/s t = 500/200 t-.5 = 500/200

OpenStudy (amistre64):

close to that ; except that last 200 should be a variable

OpenStudy (amistre64):

(5/2) - 1/2 = 500/s1 s1 = 500//2 s1 = 250 if i did it right

OpenStudy (amistre64):

50 tailwind then

OpenStudy (amistre64):

well; teh wind should slow one plave and speed the other really